OpenStudy (anonymous):

Cheryl has 16 cupcakes made. She can make 48 cupcakes an hour. A customer asks for 160 cupcakes. How many hours will it take Cheryl to make enough cupcakes for the customer?

OpenStudy (flufflepuffle):

3.333

OpenStudy (anonymous):

Okay, she a customer asks for 160 cupcakes. She already has 16 made, now how many does she have to make? divide that answer by the number of cupcakes she can make in an hour and BOOM you have your answer

OpenStudy (anonymous):

So 160 - 16 = _____ and ______/48

OpenStudy (anonymous):

Thanks so much I am clueless

OpenStudy (texaschic101):

another way is to set it up as a proportion.. 48/1 = 160/x ---- (48 cupcakes to 1 hr = 160 cupcakes to x hrs) cross multiply 48x = 160 -- divide both sides by 48 x = 160/48 x = 3.33 hrs

OpenStudy (anonymous):

actually the answer is 3 not 3.33

OpenStudy (texaschic101):

oops..that is right...because she already had 16 made,,,,so she needed 160 - 16 = 144 cupcakes. You would then set your proportion up like this : 48/1 = 144/x cross multiply 48x = 144 x = 144/48 x = 3 kmac is correct